# BF-STVSR: B-Splines and Fourier—Best Friends for High Fidelity Spatial-Temporal Video Super-Resolution

### v1

@@ 3.3 Spatial Fourier Mapper
Even with the robust motion modeling provided by the B-spline Mapper, the quality of the interpolated feature $F_{t}^{H}$ depends significantly on the features propagated from $F_{0}^{H}$ and $F_{1}^{H}$. VideoINR [7] and MoTIF [6] rely on simple MLPs to interpolate the latent features $F_{0}^{L}$ and $F_{1}^{L}$. However, implicit neural functions often struggle with capturing high-frequency details, leading to poor quality in the interpolated features, as noted in several studies [34, 41, 27]. To address this issue, LTE [22] demonstrated that using Fourier bases for spatial feature modeling significantly improves performance in arbitrary-scale super-resolution by effectively capturing dominant frequencies. Inspired by this approach, we integrate a similar strategy into our Fourier Mapper. The detail process is illustrated in Figure 2 (c). The Fourier Mapper $g_{\phi}$ predicts the dominant frequency and its amplitude of the Fourier bases for spatial features:
@@
Here, $A_{r}=g_{a}(z_{r})$ and $F_{r}=g_{f}(z_{r})$. Same as B-spline Mapper, $z_{r}=F_{t_{r}}^{L}(q_{r})$ is the nearest latent feature vector from the query coordinates $q=(x,y)$ and $\delta_{r}(=q-q_{r})$ is the relative coordinates in spatial domain. The $g_{a}$ and $g_{f}$ are the amplitude estimator ($\mathbb{R}^{C}\mapsto\mathbb{R}^{2C}$) and the frequency estimator ($\mathbb{R}^{C}\mapsto\mathbb{R}^{2C}$), respectively. By predicting dominant frequencies of query points in latent space, Fourier Mapper improves the frequency details of the interpolated features $\hat{F}_{0}^{H}$ and $\hat{F}_{1}^{H}$. After all, by linearly projecting it to $F_{0}^{H}$ and $F_{1}^{H}$ using $f_{\theta_{f}}$, which in turn enhances the quality of $F_{t}^{H}$. Note that, unlike LTE, which interpolates the estimated amplitude and frequency from low-resolution domain, our method estimates each coefficients in high-resolution domain and does not estimate the phase.
@@ 3.4 Training Objective
We follow [6] to train our model end-to-end :
$\mathcal{L}=\mathcal{L}_{\text{char}}(\hat{I}_{t}^{H},I_{t}^{H})+\beta\sum_{i=0}^{1}\mathcal{L}_{\text{char}}(\hat{M}_{i\rightarrow t}^{H},M_{i\rightarrow t}^{H}),$ (5)
where $\mathcal{L}_{\text{char}}$ is Charbonnier loss and $\beta$ is a hyper-parameter which set as 0.01. The $\hat{I}_{t}^{H}$ is the ground-truth high-resolution frame at time $t$ and $\hat{M}_{i\rightarrow t}^{H}$ is the predicted optical flow from the RAFT [42]. Note that we use RAFT only to guide our B-spline Mapper during training.
Figure 3: Qualitative comparison on arbitrary scale temporal interpolation. “Overlap” refers to the averaged image of two input frames ($t=0,1$), and the following images are interpolation results at $t=(0,1)$. (a) shows the interpolated results on in-distribution time scale (×$\times 8$), used during training. (b) shows the out-of-distribution time scale (×$\times 6$), not seen during training.
Table 3: Comparison of the degradation gap as the scale increases with various arbitrary scales in both axes. We measure the performance drop when the temporal scale is increased compared to the temporal scale ×$\times\text{6}$ at the corresponding spatial scale. Results are evaluated using PSNR (dB) and SSIM metrics. The larger (-) value means greater performance degradation.
Spatial scale $\times$4 $\times$6 $\times$12
Temporal scale $\times$6 $\Delta(\times\text{12})$ $\Delta(\times\text{16})$$\times$6 $\Delta(\times\text{12})$ $\Delta(\times\text{16})$$\times$6 $\Delta(\times\text{12})$ $\Delta(\times\text{16})$
MoTIF [6] 31.56 / 0.906 -3.79 / -0.083 -5.58 / -0.130 29.36 / 0.850 -2.58 / -0.059 -4.02 / -0.097 25.81 / 0.733 -1.09 / -0.022 -1.93 / -0.040
Ours 31.68 / 0.908 -3.62 / -0.079 -5.28 / -0.123 29.44 / 0.851 -2.38 / -0.055 -3.63 / -0.089 25.78 / 0.728 -0.91 / -0.018 -1.56 / -0.033
@@ 4.1 Experiments Setup
To our knowledge, VideoINR [7] and MoTIF [6] are the only models addressing C-STVSR. Since both models employ identical training and testing scheme, we follow the same approach unless otherwise noted. We adopt the same two-stage training strategy: for the first 450,000 iterations, the spatial scaling factor is fixed as 4, while for the remaining 150,000 iterations, it is uniformly sampled from $[1,4]$. We use the Adam optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and apply cosine annealing to decay the learning rate from $10^{-4}$ to $10^{-7}$ for every 150,000 iterations. ZoomingSlowMo [47] is used as the encoder, with a batch size of 32, and random rotation and horizontal-flipping for data augmentation. To ensure training stability, we substitute the predicted forward motion with the ground-truth forward motion with a certain probability, starting from 1.0 and gradually reducing to 0 over the first 150,000 iterations. For B-spline Mapper, we use the three-layer SIRENs as the coefficient and knot estimators, and a single fully connected layer as the dilation estimator. In Fourier Mapper, we use three-layer SIRENs as the amplitude and frequency estimators, followed by a 3$\times$3 convolution for spatial encoding. Both B-spline Mapper and Fourier Mapper have hidden dimensions of 64, with SIREN layer dimensions set to 64, 64, and 256.
@@
We use the Adobe240 dataset [40] for training, which consists of 133 videos in 720P taken by hand-held cameras. We split these videos into 100 for training, 16 for validation, and 17 for test. During training, nine sequential frames are selected from the video and the $1^{st}$ and $9^{th}$ frames are used as input reference frames. Three frames are then randomly sampled between them and used as the target ground-truth frames. For evaluation, we use Vid4[25], Adobe240 [39], and Gopro [29] datasets. Unless otherwise specified, the default spatial scale is 4, while the temporal scaling factor varies across datasets. For Vid4, temporal scale is set to $\times$2 corresponding to single frame interpolation. For Adobe240-Average and Gopro-Average, the temporal scale is set as $\times$8, representing multi-frame interpolation. Additionally, for Adobe240-center and Gopro-Center, evaluation is performed only on the center frames (i.e., the $1^{st}$, $4^{th}$, $9^{th}$ frames).
@@
We categorize baeline models into two types—continuous and fixed-scale—and conduct comparisons within each category. Here, Fixed-scale Space-Time Video Super-Resolution (Fixed-STVSR) are limited to super-resolving at fixed scaling factors in both axes that are learned during the training. First, we select two-stage Fixed-STVSR methods that combine fixed video super-resolution models (e.g., Bicubic Interpolation, EDVR [46], BasicVSR [3]) with video frame interpolation models (e.g., SuperSloMo [17], QVI [50], DAIN [1]). Second, we select one-stage Fixed-STVSR method, specifically ZoomingSlowMo [47]. For continuous methods, we select two-stage C-STVSR methods that combine continuous image super-resolution models (e.g., LIIF [5]) with video frame interpolation models (e.g., SuperSloMo [17], DAIN [1]). Lastly, we select one-stage C-STVSR methods, including TMNet [49], VideoINR [7], and MoTIF[6]. Note that TMNet is limited to $\times$4 spatial super-resolution.
@@
Metrics
We use Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) on the Y channel to evaluate the model performance.
4.2 Quantitative results
We compare our model with Fixed-STVSR methods in Table 1. For single-frame interpolation tasks in STVSR, including Vid4, GoPro-Center, and Adobe-Center, our model achieves the best performance on all datasets except Vid4. On Vid4, TMNet outperforms other models, likely due to its training on the Vimeo90K dataset [52], which has characteristics similar to Vid4. For multi-frame interpolation tasks in STVSR, represented by GoPro-Average and Adobe-Average, our model surpasses the performance of the state-of-the-art MoTIF model, which uses a pre-trained optical flow network [42] to generate temporal features during training. This improvement suggests that the B-spline Mapper and Fourier Mapper components in our model provide more robust temporal and spatial feature representations.
@@ 4.2 Quantitative results
Additionally, we compare our model with C-STVSR methods in Table 2, focusing on arbitrary-scales that were not seen during training. We perform experiments on GoPro dataset. Note that while TMNet supports continuous frame interpolation, it is limited to ×$\times\text{4}$ spatial super-resolution. Our results show that BF-STVSR achieves the best performance across all test cases, except for a minor disparity (0.03 dB lower PSNR) at $\times$6 temporal scale and $\times$6 spatial scale. This finding suggests that our B-spline Mapper effectively handles temporal interpolation even at distant time intervals. Based on Table 2, we present the performance degradation in PSNR and SSIM as the temporal scale increases from $\times$6 to $\times$12 and $\times$16 across various spatial scales in Table 3. Our results demonstrate that MoTIF experiences a more pronounced performance drop than our model under challenging conditions. Specifically, at $\times$12 spatial scale, as the temporal scale increases from $\times$6 to $\times$12, MoTIF shows a degradation of approximately -4.02 dB, while our model degrades by around -3.63 dB, indicating an improvement of nearly 0.4 dB. This suggests that our method remains robust even under extreme conditions, with temporal intervals extending up to $\times$16.
Table 5: Ablation study on the cross basis function configurations. Results are evaluated using PSNR (dB) and SSIM metrics. “Ours-cross” uses Fourier Mapper as temporal basis function and B-spline Mapper as spatial basis function. “Ours-FF” uses Fourier Mapper for both axes and “Ours-BB” uses B-spline Mapper for both axes. Bold indicates the best performance.
Vid4 Gopro-Center Gopro-Average Adobe-Center Adobe-Average
Ours 25.80 / 0.7754 31.14 / 0.8893 30.20 / 0.8799 30.84 / 0.8877 30.14 / 0.8808
Ours-cross 25.78 / 0.7747 31.10 / 0.8892 30.17 / 0.8796 30.80 / 0.8876 30.10 / 0.8804
Ours-FF 25.79 / 0.7750 31.12 / 0.8894 30.18 / 0.8798 30.84 / 0.8879 30.13 / 0.8809
Ours-BB 25.82 / 0.7753 31.07 / 0.8885 30.16 / 0.8793 30.77 / 0.8865 30.09 / 0.8797
4.3 Qualitative results
In Figure 3, qualitative results are shown, comparing our model with VideoINR and MoTIF. The results include interpolated frames for an in-distribution temporal scale ($\times$8), used during training (left), and an out-of-distribution temporal scale ($\times$6), unseen during training (right). For the in-distribution scale, BF-STVSR captures high-frequency details more effectively, particularly in the horse’s hooves and the striped shape of the handrails. For the out-of-distribution scale, BF-STVSR demonstrates superior performance in dynamic motion scenes, accurately interpolating edges of the text and the man’s face, where other methods produce blurry or ghosted frames. These results highlight our model’s ability to perform natural motion interpolation for moving objects while effectively preserving high-frequency details. Additionally, Figure 4 shows interpolated results at an extreme scale with a spatial scale of $\times$4 and a temporal scale of $\times$12. We include interpolated frames at sampled time points ($t=0.25,0.5,0.75$) along with residual intensity maps compared to ground truth frames. Our method produces sharper and more accurate results than MoTIF, especially in areas like the tire and the region next to the car window. Additional results for all time coordinates are provided in the supplementary document.
4.4 Ablation Study
Optical Flow and Position Embeddings
@@ 4.4 Ablation Study
Table 4 compares model performance with and without pre-trained optical flow network, RAFT [42], across different combinations of our proposed B-spline Mapper and Fourier Mapper modules. The first row shows the basic MoTIF [6] configuration. As seen in the second and third row, including the optical flow network with the proposed modules degrades performance. In contrast, directly using the proposed modules to extract spatial and temporal features, without the optical flow network, improves performance across all cases (last three rows). We attribute this improvement to the proposed modules facilitating the model’s effective extraction and utilization of the rich information embedded within the video, enhancing its capacity to capture complex spatial and temporal features. In addition, the fourth and fifth row in Table 4 represent the impact of the B-spline Mapper and Fourier Mapper, respectively. The fourth row is the result of using only Fourier Mapper for spatial feature encoding at BF-STVSR (last row) simply mapping the counterpart axis by concatenating latent features with target coordinates. which termed as Ours-F. The fifth row is the result of using only B-spline Mapper for temporal feature encoding at BF-STVSR (last row), which termed as Ours-B. As shown in the table, performance decreases when each mapper is used independently, but the best results are achieved when both mappers are integrated.
Different basis functions
We conduct additional experiments by varying the basis functions used for both axes (Table 5). In all configurations, the basis function model the spatial and temporal axes, while other components (e.g., linear projection for intermediate motion vectors and reliability maps) remain the same. Although the performance differences among these configurations are minimal (around 0.04 dB), we adopt the configuration using B-spline for temporal representation and Fourier for spatial representation, as it achieves the highest performance.
@@
While our method demonstrates performance improvements, there still remain certain limitations. As shown in Figure 5, existing C-STVSR models, including ours, still struggle with handling large motion. Moreover, the training process of C-STVSR models is time-consuming and computationally expensive. Addressing these challenges remains an open problem for future research.

### v2

@@ 3.3 Spatial Fourier Mapper
Even with the robust motion modeling provided by the B-spline Mapper, the quality of the interpolated feature $F_{t}^{H}$ depends significantly on the features propagated from $F_{0}^{H}$ and $F_{1}^{H}$. VideoINR [7] and MoTIF [6] rely on simple MLPs to interpolate the latent features $F_{0}^{L}$ and $F_{1}^{L}$. However, implicit neural functions often struggle with capturing high-frequency details, leading to poor quality in the interpolated features, as noted in several studies [30, 36, 23]. To address this issue, LTE [16] demonstrated that using Fourier bases for spatial feature modeling significantly improves performance in arbitrary-scale super-resolution by effectively capturing dominant frequencies. Inspired by this approach, we integrate a similar strategy into our Fourier Mapper. The detail process is illustrated in Fig 2 (c). The Fourier Mapper $g_{\phi}$ predicts the dominant frequencies and their amplitude of the Fourier bases for spatial features:
@@
Here, $A_{r}=g_{a}(z_{r})$ and $F_{r}=g_{f}(z_{r})$. Same as B-spline Mapper, $z_{r}=F_{t_{r}}^{L}(q_{r})$ is the nearest latent feature vector from the query coordinates $q=(x,y)$ and $\delta_{r}(=q-q_{r})$ is the relative coordinate in spatial domain. The $g_{a}$ and $g_{f}$ are the amplitude estimator ($\mathbb{R}^{C}\mapsto\mathbb{R}^{2C}$) and the frequency estimator ($\mathbb{R}^{C}\mapsto\mathbb{R}^{2C}$), respectively. By predicting dominant frequencies of query coordinates in latent space, Fourier Mapper improves the frequency details of the interpolated features $\hat{F}_{0}^{H}$ and $\hat{F}_{1}^{H}$. An additional linear projection $f_{\theta_{f}}$ is applied to the Fourier-embedded features, yielding refined representations of $F_{0}^{H}$ and $F_{1}^{H}$, which subsequently improve the quality of $F_{t}^{H}$. Although similar to LTE [16], the proposed Fourier Mapper estimates amplitudes and frequencies from the nearest-neighbor interpolated $z_{r}$, and does not include a phase estimator.
@@ 3.4 Training Objective
MoTIF [6] incorporates the optical flow supervision, resulting in the following training objective:
$\mathcal{L}=\mathcal{L}_{\text{char}}(\hat{I}_{t}^{H},I_{t}^{H})+\underbrace{\lambda\sum_{i=0}^{1}\mathcal{L}_{\text{char}}(\hat{M}_{i\rightarrow t}^{H},M_{i\rightarrow t}^{H})}_{\mathcal{L}_{RAFT}},$ (5)
where $\mathcal{L}_{\text{char}}$ is the Charbonnier loss, $\hat{M}_{i\rightarrow t}^{H}$ and $M_{i\rightarrow t}^{H}$ are the RAFT-predicted and model-predicted motion vectors, respectively, $\hat{I}_{t}^{H}$ and $I_{t}^{H}$ are the ground-truth and predicted high-resolution frames at time $t$, and $\lambda$ is a hyperparameter.
In contrast, our framework simplifies the objective by removing the optical flow supervision, $\mathcal{L}_{RAFT}$:
$\mathcal{L}=\mathcal{L}_{\text{char}}(\hat{I}_{t}^{H},I_{t}^{H})$ (6)
Despite this simplification, our model effectively estimates motion, achieving performance comparable to, or even better than, models trained with the optical flow supervision, $\mathcal{L}_{RAFT}$.
Figure 3: Qualitative comparison on arbitrary scale temporal interpolation. ‘‘Overlap’’ refers to the averaged image of two input frames ($t=0,1$), and the following images are interpolated results at $t\in[0,1]$. (a) shows the interpolated results on in-distribution temporal scale (×$\times 8$), used during training. (b) shows the interpolated results on out-of-distribution temporal scale (×$\times 6$), not seen during training.
Table 3: Performance comparison on the one-stage C-STVSR baselines on GoPro and Adobe240 datasets. $\mathcal{L}_{RAFT}$ refers to the optical flow supervision. Results are evaluated using VFIPS [29], FloLPIPS [9], tOF [8], and VMAF [18] metrics. All frames are interpolated by a factor of ×$\times 4$ in the spatial axis and ×$\times 8$ in the temporal axis. Red and blue indicate the best and the second best performance, respectively.
Method GoPro Adobe
[…]
VideoINR 81.13 0.151 0.519 57.96 81.15 0.145 0.574 67.08
MoTIF 81.89 0.156 0.517 59.82 81.61 0.144 0.607 68.40
BF-STVSR + $\mathcal{L}_{RAFT}$ (Ours) 83.26 0.151 0.474 61.09 84.14 0.131 0.488 70.79
BF-STVSR (Ours) 83.01 0.151 0.480 61.06 84.04 0.132 0.498 70.82
@@ 4.1 Experiments Setup
We follow the same training scheme as [7, 6] unless otherwise noted. We adopt the same two-stage training strategy: for the first 450,000 iterations, the spatial scaling factor is fixed as 4, while for the remaining 150,000 iterations, it is uniformly sampled from $[2,4]$. The $\lambda$ is set as 0.01. We use the Adam optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and apply cosine annealing to decay the learning rate from $10^{-4}$ to $10^{-7}$ for every 150,000 iterations. ZoomingSlowMo [40] is used as the encoder, with a batch size of 32, and random rotation and horizontal-flipping for data augmentation. To ensure training stability, we substitute the predicted forward motion with the ground-truth forward motion with a certain probability, starting from 1.0 and gradually reducing to 0 over the first 150,000 iterations. For B-spline Mapper, we use the three-layer SIRENs [33] as the coefficient and knot estimators, and a single fully connected layer as the dilation estimator. In Fourier Mapper, we use three-layer SIRENs as the amplitude and frequency estimators, followed by a 3$\times$3 convolutional layer for spatial encoding. Both B-spline Mapper and Fourier Mapper have hidden dimensions of 64, with SIREN layer dimensions set to 64, 64, and 256.
@@
We use the Adobe240 dataset [35] for training, which consists of 133 videos in 720P taken by hand-held cameras. During training, nine sequential frames are selected from the video and the $1^{st}$ and $9^{th}$ frames are used as input reference frames. Three frames are then randomly sampled between them and used as the target ground-truth frames. For evaluation, we use Vid4[21], Adobe240 [34], and GoPro [25] datasets. Unless otherwise specified, the default spatial scale is 4. For Vid4, temporal scale is set to $\times$2, corresponding to the center-frame interpolation. For Adobe240-Average and GoPro-Average, the temporal scale is set as $\times$8, representing multi-frame interpolation. Additionally, for Adobe240-center and GoPro-Center, evaluation is performed only on $1^{st}$, $4^{th}$, $9^{th}$ frames, representing the center-frame interpolation.
@@
We categorize baseline models into two types---continuous and fixed-scale---and conduct comparisons within each category. Here, Fixed-scale Spatial-Temporal Video Super-Resolution (Fixed-STVSR) are limited to super-resolving at fixed scaling factors in both axes that are learned during the training. First, we select two-stage Fixed-STVSR methods that combine fixed video super-resolution models (e.g., Bicubic Interpolation, EDVR [39], BasicVSR [3]) with video frame interpolation models (e.g., SuperSloMo [14], QVI [43], DAIN [1]). Second, we select one-stage Fixed-STVSR method, specifically ZoomingSlowMo [40]. For continuous methods, we select two-stage C-STVSR methods that combine continuous image super-resolution models (e.g., LIIF [5], LTE [16]) with video frame interpolation models (e.g., RIFE [13], EMA-VFI [45]). Lastly, we select one-stage C-STVSR methods, including TMNet [42], which is limited to $\times$4 spatial super-resolution, VideoINR [7], and MoTIF [6].
Evaluation Metrics
We evaluate model performance using PSNR and SSIM on the Y channel. To assess video quality, we employ VFIPS [29] and FloLPIPS [9] that primarily designed for VFI to capture perceptual similarity. Additionally, we report tOF [8] to measure temporal consistency based on the optical flow. To further evaluate video quality, we utilize VMAF [18], a perceptual metric developed for real-world video streaming applications. We measure the average VMAF score for videos encoded at 30 fps.
4.2 Quantitative results
@@
We compare our model with Fixed-STVSR methods in Table 1. For center-frame interpolation tasks in STVSR, including Vid4, GoPro-Center, and Adobe-Center, our model achieves the best performance on all datasets except Vid4. On Vid4, TMNet outperforms other models, likely due to its training on Vimeo90K dataset [44], which shares similar characteristics with Vid4. For multi-frame interpolation tasks in STVSR, represented by GoPro-Average and Adobe-Average, our model surpasses the performance of the state-of-the-art MoTIF, which uses a pre-trained optical flow network [37] to generate temporal features during training. This improvement suggests that the B-spline Mapper and Fourier Mapper provide more robust temporal and spatial feature representations. We also evaluate our model against one-stage C-STVSR methods using video quality metrics in Table 3. Our model consistently outperforms the baselines across all metrics by a significant margin, except for the FloLPIPS on GoPro dataset. This demonstrates the superior temporal consistency and perceptual quality of the proposed method. Table 2 compares the performance of the proposed method with C-STVSR methods for out-of-distribution scales on GoPro dataset. BF-STVSR achieves the best performance across all test cases, except at a $\times$16 temporal scale and $\times$4 spatial scale. This suggests that our B-spline Mapper generalizes better to unseen time intervals and effectively handles temporal interpolation. Note that in all test cases, our model performs comparably to the one with $\mathcal{L}_{RAFT}$.
4.3 Qualitative results
Fig 3 presents qualitative results comparing our model with VideoINR and MoTIF. The results include interpolated frames for an in-distribution temporal scale ($\times$8), used during training (left), and an out-of-distribution temporal scale ($\times$6), unseen during training (right). For the in-distribution scale, BF-STVSR captures high-frequency details more effectively, particularly in the horse’s hooves and the striped shape of the handrails. For the out-of-distribution scale, BF-STVSR demonstrates superior performance in dynamic motion scenes, accurately interpolating edges of the text and the man’s face, where other methods produce blurry or ghosted frames. These results highlight our model’s ability to perform natural motion interpolation for moving objects while effectively preserving high-frequency details. Additionally, Fig 4 shows interpolated results at an extreme scale with a spatial scale of $\times$4 and a temporal scale of $\times$12. We include interpolated frames at sampled time points ($t=0.25,0.5,0.75$) along with residual intensity maps compared to ground truth frames. Our method produces sharper and more accurate results than MoTIF, especially in areas like the tire and the region next to the car window.
4.4 Computational Cost and Latency
Figure 5: Computational cost (left) and inference time (right) comparison on the spatial resolution of $1280\times 720$ with different temporal scale. All frames are spatially interpolated by a factor of ×4.
To evaluate the computational efficiency of our method, we compare the FLOPs and inference time of the baselines [7, 6] and our method across different temporal scales in Fig 5. We use the fvcore library [^1] https://github.com/facebookresearch/fvcore to measure FLOPs and benchmark average inference time over 100 iterations on an NVIDIA RTX 4090 GPU. The evaluation is conducted on the spatial resolution of $1280\times 720$, with a spatial upscaling factor of ×$\times 4$ and varying temporal scales. To efficiently evaluate the B-spline function, we implement a CUDA kernel. Our method removes additional optical flow computations, enhancing efficiency. Once predicted, the B-spline representation enables lightweight motion estimation at each time step through simple linear projection, further reducing computational overhead. As shown in Fig 5, our method consistently achieves the lowest computational cost and fastest inference across all temporal resolutions.
4.5 Optical Flow and Position Embeddings
@@
30.22 / 0.8802 30.12 / 0.8808
@@
Table 4 compares model performance with and without the pre-trained optical flow network, RAFT [37], for motion modeling and the optical flow supervision, $\mathcal{L}_{RAFT}$, across different combinations of our proposed B-spline Mapper and Fourier Mapper. The first row shows the basic MoTIF [6] configuration. As seen in the second and third row, including the optical flow network with the proposed modules degrades performance. In contrast, directly using the proposed modules to extract spatial and temporal features, without the optical flow network, improves performance across all cases (last four rows). Note that even without $\mathcal{L}_{RAFT}$, our proposed model achieves similar or better performance (last row). We attribute this improvement to the ability of the proposed modules to effectively extract and utilize the rich information embedded within the video, thereby enhancing the model’s capacity to capture complex spatial and temporal features. Additionally, as shown in the fourth and fifth rows of the table, performance decreases when each mapper is used independently, but the best results are achieved when both mappers are integrated.
@@
While our method demonstrates performance improvements, there still remain certain limitations. As shown in Fig 6, existing C-STVSR models, including ours, still struggle with handling large motion. Moreover, the training process of C-STVSR models is time-consuming and computationally expensive. Addressing these challenges is left for future work.